OpenStudy (anonymous):

The table shows a direct variation. What is the constant of variation? x y 3 3.3 6.2 6.82 8 8.8

OpenStudy (igreen):

Divide. 3.3 / 3 = ? 6.82 / 6.2 = ? 8.8 / 8 = ?

OpenStudy (igreen):

Direct variation: \(\sf y = kx\)

OpenStudy (anonymous):

their that might help

OpenStudy (igreen):

Can you divide those numbers I gave you?

OpenStudy (igreen):

They should all give you the same thing.

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

1.1 1.1 1.1

OpenStudy (igreen):

Yep, so that's the constant of variation.

OpenStudy (anonymous):

okay thank you so much

OpenStudy (igreen):

No problem.

OpenStudy (anonymous):

can you help me on more?